OpenStudy (anonymous):

Jason and Kyle both choose a number from 1 to 10 at random. What is the probability that both numbers are odd?

OpenStudy (cwrw238):

there are 5 odd numbers from 1 to 10 so P(odd) = 5/10 or 1/2 P(both odd) = 1/2 * 1/2

OpenStudy (anonymous):

how many odd numbers are between 1 and 10?

OpenStudy (anonymous):

nvm